OpenStudy (anonymous):

Which of the following inequalities is graphed below? (-1,-2) and (0,2) i figured out the slope is 4 and I have the equation but it wants it with these >,< or those with the equal to sign under it.... but idk which one it would be..

OpenStudy (anonymous):

you can easily find you y-intercept

OpenStudy (anonymous):

the answers look like this \[A. y \ge -4x + 2 B. y \le 4x + 2 C. y \ge 4x+2 D. y \le -4x+2\]

OpenStudy (anonymous):

so you can eliminate A and D. do a test point

OpenStudy (anonymous):

whats a test point again?

OpenStudy (anonymous):

like (0,0) to see if the equation is true or nah

OpenStudy (anonymous):

@Tallan have you done it?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

yerp